# Trees of Graphs as Boundaries of Hyperbolic Groups

### v1

@@ 13.1. Description of $\Theta^{V}[{\mathcal{R}}_{G}]$ the tree system of V-trees associated to $G$
In order to verify the remaining conditions (2)-(4) of Definition 11.8, we again interpret Facts 13.2 and 13.3 as concerning the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$, as in the previous paragraph. In view of this, condition (2) of Definition 11.8 follows directly from Facts 13.2(1)(2) and 13.3(1)(2), condition (3) follows from Facts 13.2(1) and 13.3(1), while condition (4) is a consequence of Facts 13.2(1)(3) and 13.3(1)(3). We omit further details, thus completing the proof of the fact that the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ is modelled on the connecting system ${\mathcal{P}}^{r}_{G}$. In view of the comments at the beginning of this section and at the beginning of this subsection, this completes also the whole proof of Theorem 10.1.
@@ 13.6. Proofs of the missing parts of Theorems A and B
Proof of the implication (3)normal-$\Rightarrow$(2) in Theorem A.
@@
Proof of the part (3)normal-$\Rightarrow$(2) in Theorem B.
Suppose that condition (3) of Theorem B holds, i.e. $G$ has no rigid factor in its Bowditch JSJ splitting. Obviously, $G$ has then no rigid cluster factor in its reduced Bowditch JSJ splitting. As a consequence, all graphs $\Gamma_{z}$ in the graphical connecting system ${\mathcal{R}}_{G}$ are some thick $\theta$-graphs. It follows then from Theorem 10.1 that $\partial G$ is a regular tree of $\theta$-graphs, as asserted in condition (2) of Theorem B, which completes the proof. ∎
@@ 14. Examples
We start this section with the explanations concerning the examples provided in the introduction (at Figures 1–4). In the later part of the section, in Subsection 14.2, we discuss some Coxeter group examples.
@@
The following result characterizes the Bowditch JSJ splittings of 1-ended hyperbolic groups in terms of the features of their associated graph of groups decompositions (see Theorem 3.22 in [7], where a more general case of relatively hyperbolic groups is addressed). Recall that a splitting $\mathcal{G}$ of some group $G$, as a graph of groups, is nontrivial if no vertex group of $\mathcal{G}$ coincides with $G$. A splitting of $G$ is relative to a family $P$ of subgroups of $G$ if each subgroup from $P$ conjugates in $G$ into some vertex group of $\mathcal{G}$.
Theorem 14.1 (F. Dahmani, N. Touikan [7]).
@@
Let $S$ be a compact surface, possibly with boundary, equipped with a hyperbolic metric (so that the boundary is totally geodesic), and let $L$ be a finite collection of closed geodesics contained in the interior of $S$. Suppose that the family $L$ is filling, i.e. any connected component obtained by cutting $S$ along the union of $L$ is either a disk or an annulus, one of whose boundary components coincides with some boundary component of $S$. Let $G=\pi_{1}S$, and let $P$ be the family of (the conjugacy classes of) the cyclic subgroups of $G$ corresponding to the curves in $L$ and to the boundary components of $S$. Then $(G,P)$ is algebraically rigid, i.e. $G$ has no nontrivial splitting relative to $P$ over a finite or 2-ended subgroup.
[…]
N: Ok, I’m pretty sure this should be Theorem III.2.6.]Theorem III.2.6 in [17]). Recall that a splitting of a group $G$ as a graph of groups is minimal if the corresponding action of $G$ on the associated Bass-Serre tree $T$ is minimal, i.e. there is no $G$-invariant proper subtree $T_{0}\subset T$, $T_{0}\neq T$. Recall also that a simple closed curve on a compact surface is essential if it is homotopically non-trivial and not parallel to a boundary component.
@@
Note that the above lemma implies in particular a rather well known but useful fact that the fundamental group of a compact hyperbolic surface is freely indecomposable relative to its boundary subgroups, i.e. that the pair $(G,Q)$ as in the lemma is freely indecomposable.
@@
We now pass to a more detailed description of the Gromov boundary in one of the above examples, namely the Gromov boundary $\partial G_{2}$. We will do this by explicitely describing the graphical connecting system $\mathcal{R}_{G_{2}}$, as introduced in Section 10, and as appearing in the statement of Theorem 10.1. This will do the job, since this graphical connecting system uniquely determines the corresponding regular tree of graphs $\mathcal{X}(\mathcal{R}_{G_{2}})$, and by Theorem 10.1 we have $\partial G_{2}\cong\mathcal{X}(\mathcal{R}_{G_{2}})$.
@@
In this subsection we illustrate the results of the present paper by means of various examples of (hyperbolic) Coxeter groups, or more precisely right angled Coxeter groups (which we shortly call RACG’s). We assume that the reader is familiar with basic terminology and facts related to Coxeter groups, and we refer to the book [10] for an introduction.
In an earlier paper [24] the second author of the present paper has shown that for a large class of 1-ended hyperbolic RACG’s their Gromov boundaries are trees of graphs. This class of RACG’s can be described in terms of their defining graphs (or nerves) $N$ as follows:
@@
We denote the class of graphs $N$ satisfying the above two conditions by $\mathcal{C}_{JS}$, and the corresponding RACG’s associated to such graphs by $W_{N}$. The argument in [24] for showing that Gromov boundaries $\partial W_{N}$ of such groups are trees of graphs is different than the one given in the present paper. The trees of graphs appearing as boundaries $\partial W_{N}$ for $N\in\mathcal{C}_{JS}$ belong to a subclass consisting of so called reflection trees of graphs, as mentioned in Example 4.4 in this paper. Strictly speaking, it is shown in [24] that if $N\in\mathcal{C}_{JS}$ is a triangulation of a topological graph $X$ then $\partial W_{N}\cong\mathcal{X}^{r}(X)$ (the reflection tree of graphs $X$).
A wider class of examples has been studied in the paper [8] by Pallavi Dani and Anne Thomas. That paper contains a complete description of the Bowditch JSJ splittings for a family of 1-ended hyperbolic RACG’s whose defining graphs $N$ satisfy the following set of conditions (which appear in [8] under the name Standing Assumptions 1.1):
@@
We denote the class of graphs $N$ satisfying the above conditions by $\mathcal{C}_{DT}$. The class is strictly larger than $\mathcal{C}_{JS}$. In the light of the results in [8], and using Theorems A and B of the present paper (or rather their more precise version given as Theorem 10.1), one can indicate many more hyperbolic RACG’s $W_{N}$ with the boundary $\partial W_{N}$ recognized as an explicitely given tree of graphs (see Examples 14.5 and 14.6 below, Figures 5 and 6). One can also indicate examples with boundaries $\partial W_{N}$ of topological dimension 1 and not homeomorphic to any tree of graphs, even the ones for which all rigid factors are virtually free (see e.g. an example presented at Figure 7 below).
We now describe and discuss the promised examples from the class $\mathcal{C}_{DT}$. In all of them, as presented at Figures 5–7, we indicate their graph of groups decompositions related to the Bowditch JSJ splittings. These decompositions are obtained by applying the results of [8], and we do not provide the details, referring the interested readers directly to that paper. The underlying graphs of the indicated graph of groups decompositions in all of these examples are the quotients of the corresponding Bass-Serre trees by the actions of the groups, so one can easily deduce various properties of the corresponding Bowditch JSJ splittings (as well as of their reduced variants).
@@
Let $G=W_{N}$ be the RACG for the defining graph $N$ as presented at the left part of Figure 5. The right part of the same figure presents the graph of groups decomposition $\mathcal{G}$ of $G$ related to the Bowditsch’s JSJ splitting of this group. Vertices represented by circles are flexible, i.e. of type (v1), while those represented by little squares are of type (v2). The symbols inside circles and squares, as well as those which accompany the edges, are the generators of the corresponding special subgroups of $W_{N}$ appearing as vertex groups and edge groups of $\mathcal{G}$, respectively.
@@
There are no rigid vertices in $\mathcal{G}$, so there are also no rigid vertices in the corresponding Bass-Serre tree $T_{G}$. It follows then from Theorem B that the boundary $\partial G$ is a regular tree of $\theta$-graphs. Note that the defining graph $N$ clearly does not belong to the class $\mathcal{C}_{JS}$, so this examples goes beyond the class studied in an earlier mentioned paper [24]. (Obviously, it is easy to show many more examples of this sort, and another one, whose JSJ splitting contains a rigid factor, is described below in Example 14.6, see the group $G_{1}$ presented at Figure 6.) To get the explicit form of the regular tree of $\theta$-graphs homeomorphic to $\partial G$, i.e. to describe the corresponding graphical connecting system $\mathcal{R}_{G}$ as in Section 10, we need a closer analysis.
One deduces from the form of the graph of groups $\mathcal{G}$ that preimages in $T_{G}$ of the vertices $\langle A,B,8\rangle$ and $\langle A,C,7\rangle$ have degree 4, so that for any such preimage vertex $v$ there are precisely two edges issuing from $v$ that are mapped to each edge issuing from $\langle A,B,8\rangle$ or $\langle A,C,7\rangle$ (this is indicated at Figure 5 by little symbols ”×$\times 2$”). Preimages in $T_{G}$ of the vertex $\langle B,C\rangle$ have degree 3. Each of the 2-ended groups at vertices of type (v2) is non-orientable, because it is either an infinite dihedral group, or the product of an infinite dihedral group with a group of order 2. It follows that all flexible factors of the Bowditch JSJ splitting of $G$ are also non-orientable.
@@
To describe the set $\mathcal{A}$ of $E$-connections for $\Gamma$, denote by $a,b,c$ the edges of $\Gamma_{1}$, by $d,e,f,g$ the edges of $\Gamma_{2}$, and by $h,i,j,k$ the edges of $\Gamma_{3}$, and for each of those edges fix one of its orientations (so that these letters actually denote oriented edges). According to the description in Section 10, for each $\Gamma_{i}$ we fix a bijection between the set of edges of $\Gamma_{i}$ and the set of vertices in $T_{G}$ adjacent to the vertex $z\in Z$ representing this component. We have indicated our choices of these bijections by assigning letters $a,\dots,k$ to flexible vertices of the graph in the right part of Figure 5. Any set $Y$ of representatives of flexible vertices in $T_{G}$ can be canonically identified with the set of vertices represented by circles at Figure 5. It follows that the sets $E_{y}:y\in Y$ coincide with the sets $\{a\}$, $\{b\}$, $\{c,d,e,h,i\}$, $\{f,g\}$ and $\{j,k\}$. Due to non-orientability of all flexible factors, the set of $E$-connections has then the form
@@
Let $G_{1}=W_{N_{1}}$ and $G_{2}=W_{N_{2}}$ be the 1-ended hyperbolic RACG’s for the defining graphs $N_{i}$ as presented at the left parts of Figures 6 and 7, respectively. The central parts of these figures present the graph of groups decompositions $\mathcal{G}_{i}$ of these groups related to their Bowditch JSJ splittings. These decompositions have been obtained by referring to the results in [8]. Both underlying graphs $X_{i}$ of $\mathcal{G}_{i}$ contain a single vertex related to a rigid factor, namely the vertex represented by a hexagonal symbol, which corresponds to the special subgroup $H=\langle A,B,C,D\rangle<G_{i}$ in each of the two cases. It is also not hard to deduce from the form of $\mathcal{G}_{i}$’s that the corresponding rigid vertices in the Bass-Serre trees $T_{G_{i}}$ are not isolated, i.e they have some rigid neighbours at combinatorial distance 2. Thus, in both cases we have nontrivial rigid clusters, and we have rigid cluster factors which contain rigid factors as proper subgroups.
@@
Consider the graph of groups decompositions $\mathcal{G}^{r}_{i}$ of the groups $G_{i}$ related to their reduced Bowditch JSJ splittings. It is not hard to realize that they have forms as presented at the right parts of Figures 6 and 7. The rigid cluster factors correspond to the vertices indicated by the symbols of the bold hexagons. The rigid cluster factors in the group $G_{1}$ are the conjugates of the special subgroup $H_{1}=\langle A,B,C,D,3\rangle$, and it is not hard to see that they are virtually free. It follows then from Theorem A that the Gromov boundary $\partial G_{1}$ is a regular tree of 2-connected graphs. The explicit form of this tree of graphs, i.e. the corresponding graphical connecting system $\mathcal{R}_{G_{1}}$, can be extracted from the data in $\mathcal{G}^{r}_{1}$, and the least obvious, though still starightforward part of this extraction is the determination of the Whitehead graph associated to the rigid cluster factor $H_{1}$. We skip the details.

### v2

@@ 13.1. Description of $\Theta^{V}[{\mathcal{R}}_{G}]$ the tree system of V-trees associated to $G$
In order to verify the remaining conditions (2)-(4) of Definition 11.8, we again interpret Facts 13.2 and 13.3 as concerning the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$, as in the previous paragraph. In view of this, condition (2) of Definition 11.8 follows directly from Facts 13.2(1)(2A) and 13.3(1)(2A), condition (3) follows from Facts 13.2(1) and 13.3(1), while condition (4) is a consequence of Facts 13.2(1)(2B) and 13.3(1)(2B). We omit further details, thus completing the proof of the fact that the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ is modelled on the connecting system ${\mathcal{P}}^{r}_{G}$. In view of the comments at the beginning of this section and at the beginning of this subsection, this completes also the whole proof of Theorem 10.1.
@@ 13.6. Proofs of the missing parts of Theorems A and B
Proof of the implication (3$\Rightarrow$(2) in Theorem A.
@@
Proof of the part (3$\Rightarrow$(2) in Theorem B.
Suppose that condition (3) of Theorem B holds, i.e., $G$ has no rigid factor in its Bowditch JSJ splitting. Obviously, $G$ has then no rigid cluster factor in its reduced Bowditch JSJ splitting. As a consequence, all graphs $\Gamma_{z}$ in the graphical connecting system ${\mathcal{R}}_{G}$ are some thick $\theta$-graphs. It follows then from Theorem 10.1 that $\partial G$ is a regular tree of $\theta$-graphs, as asserted in condition (2) of Theorem B, which completes the proof. ∎
@@ 14. Examples
We start this section with the explanations concerning the examples provided in the introduction (in Figures 1–4). In the later part of the section, in Section 14.2, we discuss some Coxeter group examples.
@@
The following result characterizes the Bowditch JSJ splittings of 1-ended hyperbolic groups in terms of the features of their associated graph of groups decompositions (see Theorem 3.22 in [DT19], where a more general case of relatively hyperbolic groups is addressed). Recall that a splitting $\mathcal{G}$ of some group $G$, as a graph of groups, is nontrivial if no vertex group of $\mathcal{G}$ coincides with $G$. A splitting of $G$ is relative to a family $P$ of subgroups of $G$ if each subgroup from $P$ conjugates in $G$ into some vertex group of $\mathcal{G}$.
Theorem 14.1 (F. Dahmani, N. Touikan [DT19]).
@@
Let $S$ be a compact surface, possibly with boundary, equipped with a hyperbolic metric (so that the boundary is totally geodesic), and let $L$ be a finite collection of closed geodesics contained in the interior of $S$. Suppose that the family $L$ is filling, i.e., any connected component obtained by cutting $S$ along the union of $L$ is either a disk or an annulus, one of whose boundary components coincides with some boundary component of $S$. Let $G=\pi_{1}S$, and let $P$ be the family of (the conjugacy classes of) the cyclic subgroups of $G$ corresponding to the curves in $L$ and to the boundary components of $S$. Then $(G,P)$ is algebraically rigid, i.e., $G$ has no nontrivial splitting relative to $P$ over a finite or 2-ended subgroup.
In order to justify the above proposition, we will make use of the following lemma, which is essentially a restatement of Theorem III.2.6 in [MS84] (see also Lemmas 2.2 and 2.3 in [DS99]). Recall that a splitting of a group $G$ as a graph of groups is minimal if the corresponding action of $G$ on the associated Bass-Serre tree $T$ is minimal, i.e., there is no $G$-invariant proper subtree $T_{0}\subset T$, $T_{0}\neq T$. Recall also that a simple closed curve on a compact surface is essential if it is homotopically non-trivial and not parallel to a boundary component.
@@
Note that the above lemma implies in particular a rather well known but useful fact that the fundamental group of a compact hyperbolic surface is freely indecomposable relative to its boundary subgroups, i.e., that the pair $(G,Q)$ as in the lemma is freely indecomposable.
@@
We now pass to a more detailed description of the Gromov boundary in one of the above examples, namely the Gromov boundary $\partial G_{2}$. We will do this by explicitly describing the graphical connecting system $\mathcal{R}_{G_{2}}$, as introduced in Section 10, and as appearing in the statement of Theorem 10.1. This will do the job, since this graphical connecting system uniquely determines the corresponding regular tree of graphs $\mathcal{X}(\mathcal{R}_{G_{2}})$, and by Theorem 10.1 we have $\partial G_{2}\cong\mathcal{X}(\mathcal{R}_{G_{2}})$.
@@
In this subsection we illustrate the results of the present paper by means of various examples of (hyperbolic) Coxeter groups, or more precisely right angled Coxeter groups (which we shortly call RACG’s). We assume that the reader is familiar with basic terminology and facts related to Coxeter groups, and we refer to the book [Dav08] for an introduction.
In an earlier paper [Ś21] the second author of the present paper has shown that for a large class of 1-ended hyperbolic RACG’s their Gromov boundaries are trees of graphs. This class of RACG’s can be described in terms of their defining graphs (or nerves) $N$ as follows:
@@
We denote the class of graphs $N$ satisfying the above two conditions by $\mathcal{C}_{JS}$, and the corresponding RACG’s associated to such graphs by $W_{N}$. The argument in [Ś21] for showing that Gromov boundaries $\partial W_{N}$ of such groups are trees of graphs is different than the one given in the present paper. The trees of graphs appearing as boundaries $\partial W_{N}$ for $N\in\mathcal{C}_{JS}$ belong to a subclass consisting of so called reflection trees of graphs, as mentioned in Example 6.5 in this paper. Strictly speaking, it is shown in [Ś21] that if $N\in\mathcal{C}_{JS}$ is a triangulation of a topological graph $X$ then $\partial W_{N}\cong\mathcal{X}^{r}(X)$ (the reflection tree of graphs $X$).
A wider class of examples has been studied in the paper [DT17] by Pallavi Dani and Anne Thomas. That paper contains a complete description of the Bowditch JSJ splittings for a family of 1-ended hyperbolic RACG’s whose defining graphs $N$ satisfy the following set of conditions (which appear in [DT17] under the name Standing Assumptions 1.1):
@@
We denote the class of graphs $N$ satisfying the above conditions by $\mathcal{C}_{DT}$. The class is strictly larger than $\mathcal{C}_{JS}$. In the light of the results in [DT17], and using Theorems A and B of the present paper (or rather their more precise version given as Theorem 10.1), one can indicate many more hyperbolic RACG’s $W_{N}$ with the boundary $\partial W_{N}$ recognized as an explicitly given tree of graphs (see Examples 14.5 and 14.6 below, Figures 5 and 6). One can also indicate examples with boundaries $\partial W_{N}$ of topological dimension 1 and not homeomorphic to any tree of graphs, even the ones for which all rigid factors are virtually free (see e.g. an example presented in Figure 7 below).
We now describe and discuss the promised examples from the class $\mathcal{C}_{DT}$. In all of them, as presented in Figures 5–7, we indicate their graph of groups decompositions related to the Bowditch JSJ splittings. These decompositions are obtained by applying the results of [DT17], and we do not provide the details, referring the interested readers directly to that paper. The underlying graphs of the indicated graph of groups decompositions in all of these examples are the quotients of the corresponding Bass-Serre trees by the actions of the groups, so one can easily deduce various properties of the corresponding Bowditch JSJ splittings (as well as of their reduced variants).
@@
Let $G=W_{N}$ be the RACG for the defining graph $N$ as presented in the left part of Figure 5. The right part of the same figure presents the graph of groups decomposition $\mathcal{G}$ of $G$ related to the Bowditch JSJ splitting of this group. Vertices represented by circles are flexible, i.e., of type (v1), while those represented by little squares are of type (v2). The symbols inside circles and squares, as well as those which accompany the edges, are the generators of the corresponding special subgroups of $W_{N}$ appearing as vertex groups and edge groups of $\mathcal{G}$, respectively.
@@
There are no rigid vertices in $\mathcal{G}$, so there are also no rigid vertices in the corresponding Bass-Serre tree $T_{G}$. It follows then from Theorem B that the boundary $\partial G$ is a regular tree of $\theta$-graphs. Note that the defining graph $N$ clearly does not belong to the class $\mathcal{C}_{JS}$, so this examples goes beyond the class studied in an earlier mentioned paper [Ś21]. (Obviously, it is easy to show many more examples of this sort, and another one, whose JSJ splitting contains a rigid factor, is described below in Example 14.6, see the group $G_{1}$ presented in Figure 6.) To get the explicit form of the regular tree of $\theta$-graphs homeomorphic to $\partial G$, i.e., to describe the corresponding graphical connecting system $\mathcal{R}_{G}$ as in Section 10, we need a closer analysis.
One deduces from the form of the graph of groups $\mathcal{G}$ that preimages in $T_{G}$ of the vertices $\langle A,B,8\rangle$ and $\langle A,C,7\rangle$ have degree 4, so that for any such preimage vertex $v$ there are precisely two edges issuing from $v$ that are mapped to each edge issuing from $\langle A,B,8\rangle$ or $\langle A,C,7\rangle$ (this is indicated in Figure 5 by little symbols “×$\times 2$”). Preimages in $T_{G}$ of the vertex $\langle B,C\rangle$ have degree 3. Each of the 2-ended groups at vertices of type (v2) is non-orientable, because it is either an infinite dihedral group, or the product of an infinite dihedral group with a group of order 2. It follows that all flexible factors of the Bowditch JSJ splitting of $G$ are also non-orientable.
@@
To describe the set $\mathcal{A}$ of $E$-connections for $\Gamma$, denote by $a,b,c$ the edges of $\Gamma_{1}$, by $d,e,f,g$ the edges of $\Gamma_{2}$, and by $h,i,j,k$ the edges of $\Gamma_{3}$, and for each of those edges fix one of its orientations (so that these letters actually denote oriented edges). According to the description in Section 10, for each $\Gamma_{i}$ we fix a bijection between the set of edges of $\Gamma_{i}$ and the set of vertices in $T_{G}$ adjacent to the vertex $z\in Z$ representing this component. We have indicated our choices of these bijections by assigning letters $a,\dots,k$ to flexible vertices of the graph in the right part of Figure 5. Any set $Y$ of representatives of flexible vertices in $T_{G}$ can be canonically identified with the set of vertices represented by circles in Figure 5. It follows that the sets $E_{y}:y\in Y$ coincide with the sets $\{a\}$, $\{b\}$, $\{c,d,e,h,i\}$, $\{f,g\}$ and $\{j,k\}$. Due to non-orientability of all flexible factors, the set of $E$-connections has then the form
@@
Let $G_{1}=W_{N_{1}}$ and $G_{2}=W_{N_{2}}$ be the 1-ended hyperbolic RACG’s for the defining graphs $N_{i}$ as presented in the left parts of Figures 6 and 7, respectively. The central parts of these figures present the graph of groups decompositions $\mathcal{G}_{i}$ of these groups related to their Bowditch JSJ splittings. These decompositions have been obtained by referring to the results in [DT17]. Both underlying graphs $X_{i}$ of $\mathcal{G}_{i}$ contain a single vertex related to a rigid factor, namely the vertex represented by a hexagonal symbol, which corresponds to the special subgroup $H=\langle A,B,C,D\rangle<G_{i}$ in each of the two cases. It is also not hard to deduce from the form of $\mathcal{G}_{i}$’s that the corresponding rigid vertices in the Bass-Serre trees $T_{G_{i}}$ are not isolated, i.e they have some rigid neighbours at combinatorial distance 2. Thus, in both cases we have nontrivial rigid clusters, and we have rigid cluster factors which contain rigid factors as proper subgroups.
@@
Consider the graph of groups decompositions $\mathcal{G}^{r}_{i}$ of the groups $G_{i}$ related to their reduced Bowditch JSJ splittings. It is not hard to realize that they have forms as presented in the right parts of Figures 6 and 7. The rigid cluster factors correspond to the vertices indicated by the symbols of the bold hexagons. The rigid cluster factors in the group $G_{1}$ are the conjugates of the special subgroup $H_{1}=\langle A,B,C,D,3\rangle$, and it is not hard to see that they are virtually free. It follows then from Theorem A that the Gromov boundary $\partial G_{1}$ is a regular tree of 2-connected graphs. The explicit form of this tree of graphs, i.e., the corresponding graphical connecting system $\mathcal{R}_{G_{1}}$, can be extracted from the data in $\mathcal{G}^{r}_{1}$, and the least obvious, though still starightforward part of this extraction is the determination of the Whitehead graph associated to the rigid cluster factor $H_{1}$. We skip the details.